Say you have a spinner when with 10 different sections (all equally sized and numbered 1-10). If you plan to spin the wheel 20 times and will count the number 3 as a success.

a) What is the probability of getting the number 3 in one spin (probability of Success)?

b) Assuming this fits the criteria for a binomial distribution, what would be the expected value (mean) be?

c) What would the standard deviation be?
Answer #1

a) P(getting number 3) = 1/10 = 0.1

b) n = 20

p = 0.1

Expected value = n * p = 20 * 0.1 = 2

c) Standard deviation = sqrt(n * p * (1 - p)) = sqrt(20 * 0.1 * 0.9) = 1.34